# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

### v2

@@ 6 Conclusion
Limitations. One of the limitations of this method, like most other persistent homology based methods, is its scalability. With the increasing size of the bifiltrations, the number of sampled center points to compute Zz-Gril needs to increase. Consequently, the number of zigzag persistence computations and the length of the zigzag filtrations over which they need to be performed increase. However, recent advances in persistent homology computations in the context of ML Nigmetov & Morozov (2024) give us the hope.
@@ 7 Acknowledgement
Appendix A Missing Definitions and Proofs
@@
The $u$-shift functor $(-)_{u}:\mathbf{vec}^{\mathbb{R}^{n}}\to\mathbf{vec}^{\mathbb{R}^{n}}$, for $u\in\mathbb{R}^{n}$, is defined as follows:
@@
    Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $F:M\to N$ be a morphism. Then, the corresponding morphism $F_{u}:M_{u}\to N_{u}$ is defnied as $F_{u}(x)=F(x+u):M_{u}(x)\to N_{u}(x)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $\epsilon\in[0,\infty)$ be given. We will denote $(-)_{\bm{\epsilon}}$ to be the shift functor corresponding to the vector $\bm{\epsilon}=\epsilon(1,1,\ldots,1)$. We say $M$ and $N$ are $\epsilon$-interleaved if there are natural transformations $F:M\to N_{\bm{\epsilon}}$ and $G:N\to M_{\bm{\epsilon}}$ such that
@@
where $\varphi_{M}^{u}:M\to M_{u}$ is the natural transformation whose restriction to each $M(x)$ is the linear map $M(x\leq x+u)$ for all $x\in\mathbb{R}^{n}$.
@@
[…]
@@
Let $P$ and $Q$ be two posets. Let $F:P\to Q$ be a functor. Let $P[F\leq q]$ denote the set $P[F\leq q]\coloneqq\{p\in P:F(p)\leq q\}$. Given a persistence module $M:P\to\mathbf{vec}$, the left Kan extension of $M$ along $F$ is a functor $\text{Lan}_{F}(M):Q\to\mathbf{vec}$ given by
@@
Let $\mathbb{U}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the poset $\mathbb{U}\coloneqq\{(a,b):a\leq b\}$. Let $\mathfrak{i}:\mathbb{ZZ}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the inclusion. Let $(-)|_{\mathbb{U}}:\mathbf{vec}^{\mathbb{R}^{\mathsf{op}}\times\mathbb{R}}\to\mathbf{vec}^{\mathbb{U}}$ denote the restriction. Then, the block extension functor $E:\mathbf{vec}^{\mathbb{ZZ}}\to\mathbf{vec}^{\mathbb{U}}$ is defined as
@@
We extend this definition of interleaving distance to quasi zigzag persistence modules. Let $\mathbb{UU}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ be the analog of $\mathbb{U}$, i.e, $\mathbb{UU}\coloneqq\{(a,b,c):a\leq b\}$. Let $\tilde{\mathfrak{i}}:\mathbb{ZZ}\times\mathbb{Z}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}\times\mathbb{R}$ denote the inclusion. Then, define $\tilde{E}\coloneqq(-)|_{\mathbb{UU}}\circ\text{Lan}_{\tilde{\mathfrak{i}}}(\circ)$, analogous to the block extension functor.
@@
Let $M$ be a p.f.d quasi zigzag persistence module. If $M\cong\bigoplus_{k\in K}I_{k}$, then $\tilde{E}(M)\cong\bigoplus_{k\in K}\tilde{E}(I_{k})$, where $K$ is an indexing set.
@@
Let $I$ denote a subposet in $\mathbb{ZZ}\times\mathbb{Z}$. For ϵ+italic-$\epsilon\in\mathbb{Z}_{+}$, the $\epsilon$-thickening of $I$ is defined as
$I^{\epsilon}\coloneqq\left\{\mathbf{r}\in\mathbb{ZZ}\times\mathbb{Z}:\exists\mathbf{q}\in I\text{ such that }||\mathbf{r}-\mathbf{q}||_{\infty}\leq\epsilon\right\}.$
Recall that $(\mathbb{ZZ}\times\mathbb{Z})$ denotes the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals.
@@
Let $M$ and $N$ be two quasi zigzag persistence modules. The erosion distance is defined as:
$d_{\mathcal{E}}(M,N)\coloneqq\inf\{\epsilon\geq 0:\forall I\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\mathsf{rk}^{M}(I)\geq\mathsf{rk}^{N}(I^{\epsilon})\text{ and }\mathsf{rk}^{N}(I)\geq\mathsf{rk}^{M}(I^{\epsilon})\}.$

### v3

@@ 7 Acknowledgement
Appendix A Formal definitions and proofs for stability of Zz-Gril
@@
The $u$-shift functor $(-)_{u}\colon\mathbf{vec}^{\mathbb{R}^{n}}\to\mathbf{vec}^{\mathbb{R}^{n}}$, for $u\in\mathbb{R}^{n}$, is defined as follows:
@@
    Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $F\colon M\to N$ be a morphism. Then, the corresponding morphism $F_{u}\colon M_{u}\to N_{u}$ is defnied as $F_{u}(x)=F(x+u)\colon M_{u}(x)\to N_{u}(x)$ for all $x\in\mathbb{R}^{n}$.
@@
Let $M,N\in\mathbf{vec}^{\mathbb{R}^{n}}$. Let $\epsilon\in[0,\infty)$ be given. We will denote $(-)_{\bm{\epsilon}}$ to be the shift functor corresponding to the vector $\bm{\epsilon}=\epsilon(1,1,\ldots,1)$. We say $M$ and $N$ are $\epsilon$-interleaved if there are natural transformations $F\colon M\to N_{\bm{\epsilon}}$ and $G\colon N\to M_{\bm{\epsilon}}$ such that
@@
where $\varphi_{M}^{u}\colon M\to M_{u}$ is the natural transformation whose restriction to each $M(x)$ is the linear map $M(x\leq x+u)$ for all $x\in\mathbb{R}^{n}$.
@@
$d_{\mathcal{I}}(M,N)\coloneqq\inf\{\epsilon\geq 0\colon M\text{ and }N\text{ are }\epsilon\text{-interleaved}\}$
@@
Let $P$ and $Q$ be two posets. Let $F\colon P\to Q$ be a functor. Let $P[F\leq q]$ denote the set $P[F\leq q]\coloneqq\{p\in P\colon F(p)\leq q\}$. Given a persistence module $M\colon P\to\mathbf{vec}$, the left Kan extension of $M$ along $F$ is a functor $\text{Lan}_{F}(M)\colon Q\to\mathbf{vec}$ given by
@@
Let $\mathbb{U}\subset\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the poset $\mathbb{U}\coloneqq\{(a,b)\colon a\leq b\}$. Let $\mathfrak{i}\colon\mathbb{ZZ}\to\mathbb{R}^{\mathsf{op}}\times\mathbb{R}$ denote the inclusion. Let $(-)|_{\mathbb{U}}\colon\mathbf{vec}^{\mathbb{R}^{\mathsf{op}}\times\mathbb{R}}\to\mathbf{vec}^{\mathbb{U}}$ denote the restriction. Then, the block extension functor $E\colon\mathbf{vec}^{\mathbb{ZZ}}\to\mathbf{vec}^{\mathbb{U}}$ is defined as
@@
[…]
@@
Let $M$ be a quasi zigzag persistence module. If $M\cong\bigoplus_{k\in K}I_{k}$, then $\tilde{E}(M)\cong\bigoplus_{k\in K}\tilde{E}(I_{k})$, where $K$ is an indexing set.
@@
Let I(×)($\mathbb{ZZ}\times\mathbb{Z}$) denote the collection of all subposets in $\mathbb{ZZ}\times\mathbb{Z}$ such that their corresponding subposets in $\mathbb{Z}^{2}$ are intervals. Let $M$ and $N$ be two quasi zigzag persistence modules. The erosion distance is defined as:
$\begin{split}d_{\mathcal{E}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall I\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}(I)\geq\mathsf{rk}^{N}(I^{\epsilon})\text{ and }\\ &\mathsf{rk}^{N}(I)\geq\mathsf{rk}^{M}(I^{\epsilon})\}\end{split}$
Definition A.11.
Let $\mathcal{L}$ denote the collection of all worms in $\mathbb{ZZ}\times\mathbb{Z}$. Let $M$ and $N$ be quasi zigzag persistence modules. The erosion distance can be defined as:
$\begin{split}d_{\mathcal{E}}^{\mathcal{L}}(M,N)\coloneqq\inf\limits_{\epsilon\geq 0}\{&\forall\boxed{\mathbf{p}}^{2}_{\delta}\in\textbf{I}(\mathbb{ZZ}\times\mathbb{Z}),\\ &\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\text{ and }\\ &\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)\}.\end{split}$
@@
Proposition 3.6.
Proof.
First, it is obvious that $d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{E}}(M,N)$. Now, $d_{\mathcal{I}}(M,N)\coloneqq d_{\mathcal{I}}(\tilde{E}(M),\tilde{E}(N))$. Let $I$ be a subposet in $\mathbb{ZZ}\times\mathbb{Z}$ such that its corresponding subposet in $\mathbb{Z}^{2}$ is an interval. By Lemma A.8 and Lemma A.9, and the fact that generalized rank over a given subposet counts the number of intervals that contain the given subposet, we get $\mathsf{rk}^{M}(I)=\mathsf{rk}^{\tilde{E}(M)}(\tilde{\mathfrak{i}}(I)\cap\mathbb{UU})$. This gives us $d_{\mathcal{E}}(M,N)=d_{\mathcal{E}}(\tilde{E}(M),\tilde{E}(N))$. In Kim & Mémoli (2021), the authors show that $d_{\mathcal{E}}(V,W)\leq d_{\mathcal{I}}(V,W)$, where $V,W\colon\mathbb{R}^{n}\to\textbf{vec}$ are $\mathbb{R}^{n}$-indexed persistence modules. Thus, we get $d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{E}}(M,N)=d_{\mathcal{E}}(\tilde{E}(M),\tilde{E}(N))\leq d_{\mathcal{I}}(\tilde{E}(M),\tilde{E}(N))=d_{\mathcal{I}}(M,N)$.
Theorem 3.7.
Let $M$ and $N$ be two quasi zigzag persistence modules. Let $\lambda^{M}$ and $\lambda^{N}$ denote the Zz-Gril functions of $M$ and $N$ respectively. Then,
$\lambda^{M}-\lambda^{N}||_{\infty}=d_{\mathcal{E}}^{\mathcal{L}}(M,N)\leq d_{\mathcal{I}}(M,N).$
Theorem 3.7.
Proof.
We show that $\lambda^{M}-\lambda^{N}||_{\infty}=d_{\mathcal{E}}^{\mathcal{L}}(M,N)$.
To see $\lambda^{M}-\lambda^{N}||_{\infty}\leq d_{\mathcal{E}}^{\mathcal{L}}(M,N)$, fix $\mathbf{p},k$, and let $\lambda^{M}(\mathbf{p},k)=\delta_{M}$ and $\lambda^{N}(\mathbf{p},k)=\delta_{N}$. WLOG, assume $\delta_{M}\geq\delta_{N}$. Let $d_{\mathcal{E}}^{\mathcal{L}}(M,N)=\epsilon$. Therefore, by definition of $d_{\mathcal{E}}^{\mathcal{L}}$, we have $\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta_{M}}\right)=k$ and $\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta_{N}+\epsilon}\right)\leq\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta_{N}}\right)=k$. Thus, by the definition of Zz-Gril, we get $\delta_{N}+\epsilon\geq\delta_{M}$, i.e., $\delta_{M}-\delta_{N}\leq\epsilon=d^{\mathcal{L}}_{\mathcal{E}}$.
To see $\lambda^{M}-\lambda^{N}||_{\infty}\geq d_{\mathcal{E}}^{\mathcal{L}}(M,N)$, fix $\mathbf{p},k$, and let $|\lambda^{M}(\mathbf{p},k)-\lambda^{N}(\mathbf{p},k)|=\epsilon$. Let $\boxed{\mathbf{p}}^{2}_{\delta}$ be a worm. Let $k=\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)$. Then, $\lambda^{N}(\mathbf{p},k)\geq\delta+\epsilon$. Thus, $|\lambda^{M}(\mathbf{p},k)-\lambda^{N}(\mathbf{p},k)|=\epsilon$ and $\lambda^{N}(\mathbf{p},k)\geq\delta+\epsilon$ give $\lambda^{M}(\mathbf{p},k)\geq d$. Therefore, $\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq k=\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)$. Similarly, we can show $\mathsf{rk}^{N}\left(\boxed{\mathbf{p}}^{2}_{\delta}\right)\geq\mathsf{rk}^{M}\left(\boxed{\mathbf{p}}^{2}_{\delta+\epsilon}\right)$. Thus, by definition of $d_{\mathcal{E}}^{\mathcal{L}}$, we get that $\lambda^{M}-\lambda^{N}||_{\infty}\geq d_{\mathcal{E}}^{\mathcal{L}}(M,N)$.
Appendix B Experimental Details
Here, we report additional details about our experiments. We begin by including a detailed description of the UEA datasets in Table 8. Further, we have a comparison of treating a multivariate time series as a sequence of point clouds versus treating it as a sequence of graphs as the input to the Zz-Gril framework. We report the results of this experiment in Table 4. We can see that there is no clear winner. On some datasets, Zz-Gril extracts more relevant information from sequences of point clouds while on some datasets, information from sequences of graphs performs better. We have a visualization of Zz-Gril in Figure 7. We can see in the figure that the Zz-Gril topological signature is different for samples of different classes and very similar for samples in the same class.
For all the experiments with sequence of point clouds, we choose a window size that is $\max(5,series\_length//128)$ and overlap $\max(4,0.7*window\_size)$. For the experiments with sequence of graphs, we choose a window size of $\min(series\_length/5,128)$ and an overlap of 0.7window_size0.7*window\_size. Then, from the complete graph, we randomly choose a percentage between 65 and 75 of the edges depending on their weights. This is to ensure that the number of edges does not remain the same for all the graphs in the sequence. For all our experiments, we use 36 center points to compute Zz-Gril at. For model parameters, mostly, we use the same parameters as specified by the respective models Liu et al. (2024); Zhao et al. (2023). We notice that the training is sensitive to learning rate and we optimize the learning rate between 1e31e-3 and 5e55e-5 for different datasets. These choice of hyperparameters are based on preliminary experiments, the results of which we report in Table 9, Table 10 , Table 11 and Table 12. The results reported here are for TodyNet+Zz-Gril model.
Dataset Type No. of series Series length No. of classes Train size Test size
FingerMovements EEG 28 50 2 316 100
Heartbeat AUDIO 61 405 2 204 205
MotorImagery EEG 64 3000 2 278 100
NATOPS HAR 24 51 6 180 180
SelfRegulationSCP2 EEG 7 1152 2 200 180
Table 8: Information about UEA Datasets used for experiments
Dataset 36 center points 25 center points 64 center points
FingerMovements 0.660 0.620 0.680
SelfRegulationSCP2 0.600 0.578 0.600
Table 9: Selection of number of center points to compute Zz-Gril
Dataset $\mathbf{max(0.7*window\_size,5)}$ $\mathbf{max(0.3*window\_size,2)}$ $\mathbf{max(0.5*window\_size,3)}$
FingerMovements 0.660 0.620 0.640
SelfRegulationSCP2 0.600 0.550 0.578
Table 10: Selection of overlap between time windows for time-series modelled as a sequence of graphs. The window_sizewindow\_size is fixed to be max(128//series_len,5max(128//series\_len,5).
Dataset $\mathbf{max(series\_len//128,5)}$ $\mathbf{max(series\_len//64,10)}$ $\mathbf{max(series\_len//256,3)}$
FingerMovements 0.660 0.640 0.580
SelfRegulationSCP2 0.600 0.578 0.550
Table 11: Selection of window size for time-series modelled as a sequence of graphs. The overlap is fixed to be max(4,0.7window_size)max(4,0.7*window\_size).
Dataset Threshold (45% - 55%) Threshold (65% - 75%) Threshold (85% - 95 %)
FingerMovements 0.600 0.660 0.630
SelfRegulationSCP2 0.550 0.600 0.550
Table 12: Selection of thresholding value range for retaining edges in the graph.